OpenStudy (anonymous):

Which value of x makes the open sentence true? 7 – x = 3x + 5 A.x = 0.5 B.x = 1.5 C.x = 3 D.x = 2

OpenStudy (jango_in_dtown):

7-x=3x+5 imply 7-5=3x+x implies 2=4x implies x=1/2 i.e. 0.5

OpenStudy (anonymous):

its A. 0.5